# RanPAC: Random Projections and Pre-trained Models for Continual Learning

### v1

@@ Abstract
Continual learning (CL) aims to incrementally learn different tasks (such as classification) in a non-stationary data stream without forgetting old ones. Most CL works focus on tackling catastrophic forgetting under a learning-from-scratch paradigm. However, with the increasing prominence of foundation models, pre-trained models equipped with informative representations have become available for various downstream requirements. Several CL methods based on pre-trained models have been explored, either utilizing pre-extracted features directly (which makes bridging distribution gaps challenging) or incorporating adaptors (which may be subject to forgetting). In this paper, we propose a concise and effective approach for CL with pre-trained models. Given that forgetting occurs during parameter updating, we contemplate an alternative approach that exploits training-free random projectors and class-prototype accumulation, which thus bypasses the issue. Specifically, we inject a frozen Random Projection layer with nonlinear activation between the pre-trained model’s feature representations and output head, which captures interactions between features with expanded dimensionality, providing enhanced linear separability for class-prototype-based CL. We also demonstrate the importance of decorrelating the class-prototypes to reduce the distribution disparity when using pre-trained representations. These techniques prove to be effective and circumvent the problem of forgetting for both class- and domain-incremental continual learning. Compared to previous methods applied to pre-trained ViT-B/16 models, we reduce final error rates by between 10% and 62% on seven class-incremental benchmark datasets, despite not using any rehearsal memory. We conclude that the full potential of pre-trained models for simple, effective, and fast continual learning has not hitherto been fully tapped.
@@ 1 Introduction
Although forgetting still occurs in CL with such transfer-learning L2P; zhou2023revisiting; zhang2023slca (whether using conventional fine-tuning Kornblith.18 or more-recent Parameter-Efficient Transfer Learning (PETL) methods such as chen2022adaptformer; NEURIPS2022_00bb4e41; jia2022vpt), commencing CL with a powerful feature-extractor has opened up new ideas for avoiding forgetting that are unlikely to work when training from scratch. Such new methods have been applied successfully to pre-trained transformer networks pelosin2022simpler; Ermis22; L2P; DualPrompt; smith2023codaprompt; janson2023simple; zhou2023revisiting; zhang2023slca; smith2023closer; S-prompts, CNNs hayes2020lifelong; Hocquet20; mehta2021empirical; Jie_2022_CVPR; panos2023session and multi-modal vision-language models ding2022dont.
@@
    We show that random projections are particularly useful when also using PETL methods with first-session training (see Section 2.1). Accuracy with this combination approaches the CL joint training upper bound on some datasets. For ViT-B/16 pre-trained models, we report the highest-to-date rehearsal-free CL accuracies on all class-incremental and domain-incremental datasets we tested on, with large margins when compared to past CP strategies.
@@ 2.1 Three strategies for CL with strong pre-trained models
CPs are well suited to CL generally icarl; Lange21; Mai21; UCIR and for application to pre-trained models janson2023simple; zhou2023revisiting; panos2023session, because when the model from which feature vectors are extracted is frozen, CPs accumulated across $T$ tasks will be identical regardless of the ordering of the tasks. Moreover, their memory cost is low compared with using a memory buffer of rehearsal samples, the strategy integral to many CL methods wang2023comprehensive.
@@ 3.3 CP-based classification using Gram matrix inversion
We will also use incrementally calculate second-order feature statistics, but create a simplification compared with LDA, by using the Gram matrix of the features, ${\bf G}$, and ${\bf c}_{y}$ (CPs with averaging dropped), to obtain the predicted label
@@ 4.1 Why second order statistics matter for CPs
For comparison with CP approaches, we jointly trained, on all 200 classes, a linear probe softmax classifier. We treat the weights of the joint probe as class-prototypes for this exercise and then find the Pearson correlation coefficients between each pair of prototypes as shown for the first 10 classes in Fig. 2 (right). Compared with the linear probe, very high off-diagonal correlations are clearly observed when using NCM, with a mean value more than twice that of the original ImageNet-1K training data treated in the same way. This illustrates the extent of the domain shift for the downstream dataset. However, these correlations are mostly removed when using Eqn (2).
Fig. 2 (left) shows that high correlation coefficients coincide with poorly calibrated cosine similarities between class-prototypes and training samples, both for true class comparisons and inter-class comparisons. However, when using Eqn. (2), the result (third row) is to increase the training-set accuracy from 64% to 75%, coinciding with reduced overlap between inter- and true-class similarity distributions, and significantly reduced off-diagonal correlation coefficients between CPs. The net result is linear classification weights that are much closer to those produced by the jointly-trained linear probe. These results are consistent with known mathematical relationships between Eqn (2) and decorrelation, which we outline in Appendix B.4.4.
Figure 2: Left: Histograms of similarities between class-prototypes and feature vectors for the training samples of the Imagenet-R dataset, using features extracted from the ViT-B/16 transformer model pre-trained on ImageNet-1K. Right: Pearson correlation coefficients (CCs) for 10 pairs of class-prototypes.
@@ 4.2 Detailed overview and intuition for Random Projections
This assertion is supported by the bottom row of Fig. 2, which shows that the application of an RP layer of size $M=2000$ reduces the overlap between the similarity histograms further than otherwise, and shifts the in-class histogram to the right, coinciding with higher accuracy, surpassing that of the joint linear probe, consistent with our findings in Section 5. We note that these RPs can be generated once and then frozen for use in all continual learning stages, enabling a simple and efficient algorithm.
@@ 4.5 Memory usage of RP layer
The RP layer increases the total number of trainable parameters by a factor of $1+(M-L)/L$, while an additional $LM$ parameters are frozen and untrainable. For typical values of $L=768$ and $K=200$, the injection of an RP layer of size $M=10000$ therefore creates $\sim 10$ times the number of trainable parameters, and adds $\sim 10$M non-trainable parameters. Although this is a significant number of new parameters, it is still small compared with the overall size of the ViT-B/16 model, with its 84 millions parameters. Moreover, the weights of ${\bf W}$ can be bipolar instead of Gaussian and if stored using a single bit for each element contribute only a tiny fraction to additional model memory. During training only, we also require updating an $M\times M$ Gram matrix, which is smaller than the $K$ $(L\times L)$ covariance matrices of the SLCA fine-tuning approach zhang2023slca if $M<\sqrt{K}L$.
@@ 5 Experiments
Table 2: Comparison with fine-tuning strategies for CIL. Results for SLCA are directly from zhang2023slca. Joint means full fine-tuning of the entire pre-trained ViT network. Notes: L2 smith2023closer achieved 76.1% on Imagenet-R, but reported no other ViT results. Cells containing - indicates SLCA did not report results, and no codebase is available yet. Superscript $*{}^{*}$: the number of CUB training samples used by zhang2023slca is smaller than that defined by zhou2023revisiting, which we use.
@@ 6 Conclusion
Limitations: The value of Eqs (4) and  (5) are completely reliant on supply of a good generic feature extractor. For this reason, they are unlikely to be as powerful if used in CL methods that train networks from scratch. However, it is possible that existing CL methods that utilise self-supervised learning, or otherwise create good feature extractor backbones could leverage similar approaches to ours for downstream CL tasks.
@@ Appendix A Overview of RanPAC and comparison to other strategies
Figure A1 provides a graphical overview of the two phases in Algorithm 1. Table A1 provides a summary of different strategies for leveraging pre-trained models for Continal Learning (CL) and how our own method, RanPAC, compares.
Figure A1: Overview of RanPAC for CL classification. In Phase 1 of Algorithm 1 we optionally inject parameters for a Parameter-Efficient Transfer Learning (PETL) method into a frozen pre-trained model (PTM). The PETL parameters are trained only on Task 1 (the ‘first-session’) of a set of $T$ continual learning tasks, to help bridge the domain gap, as in zhou2023revisiting; panos2023session. Then in Phase 2, first, $L$-dimensional feature vectors, ${\bf f}$, are extracted from the network after completion of learning in Phase 1 (now frozen). Then, the extracted feature vectors are randomly projected to dimension $M$ (typically $M>L$) using frozen weights ${\bf W}$, followed by nonlinear activation, $\phi$, to obtain new feature vectors, ${\bf h}=\phi({\bf f}^{\top}{\bf W})$. Training in Phase 2 is comprised of continual iterative updating of class prototypes and the Gram matrix, followed by a matrix inversion to compute $M\times N$ matrix ${\bf W}_{\rm o}$, following the end of each Task’s training. These weights can be thought of as decorrelated class prototypes that linearly weight the features in ${\bf h}$ to obtain class predictions in ${\bf y}$.
@@
 L2P; DualPrompt; smith2023codaprompt; S-prompts  zhang2023slca; smith2023closer  janson2023simple; zhou2023revisiting; panos2023session RanPAC
@@ B.3 Connection to least squares
where ${\bf Y}_{\rm train}$ is a $N\times K$ one-hot encoded target matrix, and ${\bf H}$ is as defined in Section 4.3. Eqn. (12) is well known (murphy2013machine, Eqn (7.33), p. 226) to be the minimum mean squared error solution to the $l_{2}$ regularized set of equations defined by
@@ B.4.2 Relationship to LDA and Mahalanobis distance
The form of Eqn. (2) resembles Linear Discriminant Analysis (LDA) classification (murphy2013machine, Eqn. (4.38), p. 104). For LDA, the score from which the predicted class for ${\bf f}_{\rm test}$ is chosen is commonly expressed in a weighting and bias form
@@ C.2 Training details
For Phase 2 in Algorithm 1, the training data was used as follows. For each sample, features were extracted from a frozen pretrained model, in order to update the ${\bf G}$ and ${\bf C}$ matrices. We then computed ${\bf W}_{\rm o}$ using matrix inversion and multiplication. Hence, no actual training is required.
@@
For reported data using linear probes or full fine-tuning, we also used batch sizes of $48$, a learning rate of $0.01$, weight decay of $0.0005$, momentum of $0.9$ and training for $20$ epochs. However, we used a fixed learning rate schedule, with no reductions in learning rate.
@@
C.3 Compute
@@ F.3 Comparison of impact of Phase 1 for different numbers of CIL tasks
Therefore, in this section we show in Table A4 results only for the case where variability has greater effect, i.e. when Phase 1 is included, and AdaptMLP chosen. VTAB is excluded from this analysis, since it is a special case where it makes sense to consider only the case of $T=5$.
@@
Table A4: Comparison of CIL results for different number of tasks. For this table, the AdaptMLP PETL method was used. In most cases, Average Accuracy decreases as $T$ increases, with $T=20$ typically not significantly better than the case of no PETL.
@@ F.4 Task Agnostic Continual Learning
Unlike CIL and DIL, ‘task agnostic’ CL is a scenario where there is no clear concept of a ‘task’ during training zeno2019task. To illustrate the flexibility of RanPAC, we show here that it is simple to apply it to task agnostic continual learning. We use the Gaussian scheduled CIFAR100 protocol of L2P, which was adapted from Shanahan21. We use 200 ‘micro-tasks’ which sample from a gradually shifting subset of classes, with 5 batches of 48 samples in each micro-task. There are different possible choices for how to apply Algorithm 1. For instance, the ‘first session’ for Phase 1 could be defined as a particular total number of samples trained on, e.g. 10% of the anticipated total number of samples. Then in Phase 2, the outer for loop over tasks could be replaced by a loop over all batches, or removed entirely. In both cases, the result for ${\bf G}$ and ${\bf C}$ will be unaffected. The greater challenge is in determining $\lambda$, but generally for a large number of samples, $\lambda$ can be small, or zero. For a small number of training samples, a queue of samples could be maintained such that the oldest sample in the queue is used to update ${\bf G}$ and ${\bf C}$, with all newer samples used to compute $\lambda$ if inference is required, and then all samples in the buffer added to ${\bf G}$ and ${\bf C}$.

### v2

@@ Abstract
Continual learning (CL) aims to incrementally learn different tasks (such as classification) in a non-stationary data stream without forgetting old ones. Most CL works focus on tackling catastrophic forgetting under a learning-from-scratch paradigm. However, with the increasing prominence of foundation models, pre-trained models equipped with informative representations have become available for various downstream requirements. Several CL methods based on pre-trained models have been explored, either utilizing pre-extracted features directly (which makes bridging distribution gaps challenging) or incorporating adaptors (which may be subject to forgetting). In this paper, we propose a concise and effective approach for CL with pre-trained models. Given that forgetting occurs during parameter updating, we contemplate an alternative approach that exploits training-free random projectors and class-prototype accumulation, which thus bypasses the issue. Specifically, we inject a frozen Random Projection layer with nonlinear activation between the pre-trained model’s feature representations and output head, which captures interactions between features with expanded dimensionality, providing enhanced linear separability for class-prototype-based CL. We also demonstrate the importance of decorrelating the class-prototypes to reduce the distribution disparity when using pre-trained representations. These techniques prove to be effective and circumvent the problem of forgetting for both class- and domain-incremental continual learning. Compared to previous methods applied to pre-trained ViT-B/16 models, we reduce final error rates by between 10% and 62% on seven class-incremental benchmark datasets, despite not using any rehearsal memory. We conclude that the full potential of pre-trained models for simple, effective, and fast continual learning has not hitherto been fully tapped. Code is available at https://github.com/RanPAC/RanPAC.
@@ 1 Introduction
Although forgetting still occurs in CL with such transfer-learning L2P; zhou2023revisiting; zhang2023slca (whether using conventional fine-tuning Kornblith.18 or more-recent Parameter-Efficient Transfer Learning (PETL) methods such as chen2022adaptformer; NEURIPS2022_00bb4e41; jia2022vpt), commencing CL with a powerful feature-extractor has opened up new ideas for avoiding forgetting that are unlikely to work when training from scratch. Such new methods have been applied successfully to pre-trained transformer networks pelosin2022simpler; Ermis22; L2P; DualPrompt; smith2023codaprompt; janson2023simple; zhou2023revisiting; zhang2023slca; smith2023closer; S-prompts, CNNs hayes2020lifelong; Hocquet20; mehta2021empirical; Jie_2022_CVPR; Wu_2022_CVPR; panos2023session and multi-modal vision-language models ding2022dont.
@@
    We show that random projections are particularly useful when also using PETL methods with first-session training (see Section 2.1). Accuracy with this combination approaches the CL joint training upper bound on some datasets (Table 2). For ViT-B/16 pre-trained models, we report the highest-to-date rehearsal-free CL accuracies on all class-incremental and domain-incremental datasets we tested on, with large margins when compared to past CP strategies.
@@ 2.1 Three strategies for CL with strong pre-trained models
CPs are well suited to CL generally icarl; Lange21; Mai21; UCIR and for application to pre-trained models janson2023simple; zhou2023revisiting; panos2023session, because when the model from which feature vectors are extracted is frozen, CPs accumulated across $T$ tasks will be identical regardless of the ordering of the tasks. Moreover, their memory cost is low compared with using a rehearsal buffer, the strategy integral to many CL methods wang2023comprehensive.
@@ 3.3 CP-based classification using Gram matrix inversion
We will also use incrementally calculated second-order feature statistics, but create a simplification compared with LDA (see Appendix B.4), by using the Gram matrix of the features, ${\bf G}$, and ${\bf c}_{y}$ (CPs with averaging dropped), to obtain the predicted label
@@ 4.1 Why second order statistics matter for CPs
For comparison with CP approaches, we jointly trained, on all 200 classes, a linear probe softmax classifier. We treat the weights of the joint probe as class-prototypes for this exercise and then find the Pearson correlation coefficients between each pair of prototypes as shown for the first 10 classes in Fig. 2 (right). Compared with the linear probe, very high off-diagonal correlations are clearly observed when using NCM, with a mean value more than twice that of the original ImageNet-1K training data treated in the same way. This illustrates the extent of the domain shift for the downstream dataset. However, these correlations are mostly removed when using Eqn (2). Fig. 2 (left) shows that high correlation coefficients coincide with poorly calibrated cosine similarities between class-prototypes and training samples, both for true class comparisons (similarities between a sample’s feature vector and the CP for the class label corresponding to that sample.) and inter-class comparisons (similarities between a sample’s feature vector and the class prototypes for the set of N-1 classes not equal to the sample’s class label). However, when using Eqn. (2), the result (third row) is to increase the training-set accuracy from 64% to 75%, coinciding with reduced overlap between inter- and true-class similarity distributions, and significantly reduced off-diagonal correlation coefficients between CPs. The net result is linear classification weights that are much closer to those produced by the jointly-trained linear probe. These results are consistent with known mathematical relationships between Eqn (2) and decorrelation, which we outline in Appendix B.4.4.
Figure 2: Left: Histograms of similarities between class-prototypes and feature vectors from the ViT-B/16 transformer model pre-trained on ImageNet-1K, for the training samples of the Imagenet-R dataset. Right: Pearson correlation coefficients (CCs) for 10 pairs of class-prototypes. Reduced correlations between CPs of different classes (right), coincides with better class separability (left).
@@ 4.2 Detailed overview and intuition for Random Projections
This assertion is supported by the bottom row of Fig. 2, which shows that the application of an RP layer of size $M=2000$ (chosen arbitrarily, for the purpose of illustration) reduces the overlap between the similarity histograms further than otherwise, and shifts the in-class histogram to the right, coinciding with higher accuracy, surpassing that of the joint linear probe, consistent with our findings in Section 5. We note that these RPs can be generated once and then frozen for use in all continual learning stages, enabling a simple and efficient algorithm.
@@ 4.5 Memory usage of RP layer
The RP layer increases the total memory required by the class prototypes by a factor of $1+(M-L)/L$, while an additional $LM$ parameters are frozen and untrainable. For typical values of $L=768$ and $K=200$, the injection of an RP layer of size $M=10000$ therefore creates $\sim 10$ times the number of trainable parameters, and adds $\sim 10$M non-trainable parameters. Although this is a significant number of new parameters, it is still small compared with the overall size of the ViT-B/16 model, with its 84 millions parameters. Moreover, the weights of ${\bf W}$ can be bipolar instead of Gaussian and if stored using a single bit for each element contribute only a tiny fraction to additional model memory. During training only, we also require updating an $M\times M$ Gram matrix, which is smaller than the $K$ $(L\times L)$ covariance matrices of the SLCA fine-tuning approach zhang2023slca if $M<\sqrt{K}L$. L2P L2P, DualPrompt DualPrompt and ADaM zhou2023revisiting each use $\sim$0.3M–0.5M trainable parameters. For $K=200$ classes and $M=10000$, RanPAC uses a lot more ($\sim$2–2.5M, depending on the PETL method), comparable to CODA-Prompt smith2023codaprompt. RanPAC also uses 10M untrained parameters, but we highlight that these are not trainable. Moreover, $M$ need not be as high as $10000$ (Table A5).
@@ 5 Experiments
Table 2: Comparison with fine-tuning strategies for CIL. Results for SLCA are directly from zhang2023slca. ‘Joint’ means full fine-tuning of the entire pre-trained ViT network, i.e. a continual learning upper bound. Notes: L2 smith2023closer achieved 76.1% on Imagenet-R, but reported no other ViT results. Cells containing - indicates SLCA did not report results, and no codebase is available yet. Superscript $*{}^{*}$: the number of CUB training samples used by zhang2023slca is smaller than that defined by zhou2023revisiting, which we use.
@@ 6 Conclusion
Limitations: The value of Eqs (4) and  (5) are completely reliant on supply of a good generic feature extractor. For this reason, they are unlikely to be as powerful if used in CL methods that train networks from scratch. However, it is possible that existing CL methods that utilise self-supervised learning, or otherwise create good feature extractor backbones could leverage similar approaches to ours for downstream CL tasks. As discussed in Section 4.5, RanPAC uses additional parameters compared to methods like L2P. However, this is arguably worth trading-off for the simplicity of implementation and low-cost training.
@@
Acknowledgements
This work was supported by the Centre for Augmented Reasoning at the Australian Institute for Machine Learning, established by a grant from the Australian Government Department of Education. Dong Gong is the recipient of an Australian Research Council Discovery Early Career Award (project number DE230101591) funded by the Australian Government. We would like to thank Sebastien Wong of DST Group, Australia, and Lingqiao Liu of the Australian Institute for Machine Learning, The University of Adelaide, for valuable suggestions and discussion.
@@ Appendix A Overview of RanPAC and comparison to other strategies
Figure A1 provides a graphical overview of the two phases in Algorithm 1. Table A1 provides a summary of different strategies for leveraging pre-trained models for Continual Learning (CL) and how our own method, RanPAC, compares.
Figure A1: Overview of RanPAC for CL classification. In Phase 1 of Algorithm 1 we optionally inject parameters for a Parameter-Efficient Transfer Learning (PETL) method into a frozen pre-trained model (PTM). The PETL parameters are trained only on Task 1 (the ‘first-session’) of a set of $T$ continual learning tasks, to help bridge the domain gap, as in zhou2023revisiting, panos2023session. Then in Phase 2, first, $L$-dimensional feature vectors, ${\bf f}$, are extracted from the network after completion of learning in Phase 1 (now frozen). Then, the extracted feature vectors are randomly projected to dimension $M$ (typically $M>L$) using frozen weights ${\bf W}$, followed by nonlinear activation, $\phi$, to obtain new feature vectors, ${\bf h}=\phi({\bf f}^{\top}{\bf W})$. Training in Phase 2 is comprised of continual iterative updating of class prototypes and the Gram matrix, followed by a matrix inversion to compute $M\times N$ matrix ${\bf W}_{\rm o}$, following the end of each Task’s training. These weights can be thought of as decorrelated class prototypes that linearly weight the features in ${\bf h}$ to obtain class predictions in ${\bf y}$.
@@
 L2P, DualPrompt, smith2023codaprompt, S-prompts  zhang2023slca, smith2023closer  janson2023simple, zhou2023revisiting, panos2023session RanPAC
@@ B.3 Connection to least squares
where ${\bf Y}_{\rm train}$ is a $N\times K$ one-hot encoded target matrix, and ${\bf H}$ is as defined in Section 4.3. Eqn. (12) is well known [murphy2013machine, Eqn (7.33), p. 226] to be the minimum mean squared error solution to the $l_{2}$ regularized set of equations defined by
@@ B.4.2 Relationship to LDA and Mahalanobis distance
The form of Eqn. (2) resembles Linear Discriminant Analysis (LDA) classification [murphy2013machine, Eqn. (4.38), p. 104]. For LDA, the score from which the predicted class for ${\bf f}_{\rm test}$ is chosen is commonly expressed in a weighting and bias form
@@ C.2 Training details
For Phase 2 in Algorithm 1, the training data was used as follows. For each sample, features were extracted from a frozen pretrained model, in order to update the ${\bf G}$ and ${\bf C}$ matrices. We then computed ${\bf W}_{\rm o}$ using matrix inversion and multiplication. Hence, no SGD based weight updates are required in Phase 2.
@@
For reported data in Table 1 using linear probes we used batch sizes of $128$, a learning rate of $0.01$ in the classification head, weight decay of $0.0005$, momentum of $0.9$ and training for $30$ epochs. For full fine-tuning (Table 2), we used the same settings, but additionally used a learning rate in the body (the pre-trained weights of the ViT backbone) of 0.0001. We used a fixed learning rate schedule, with no reductions in learning rate. We found for fine-tuning that the learning rate lower in the body than the head was essential for best performance.
@@
C.3 Training and Inference Speed
The speed for inference with RanPAC is negligibly different to the speed of the original pre-trained network, because both the RP layer and the output linear classification head (comprised from decorrelated class prototypes) are implemented as simple fully-connected layers on top of the underlying network. For training, Phase 1 trains PETL parameters using SGD for 20 epochs, on $(1/T)$’th of the training set, so is much faster than joint training. Phase 2 is generally only slightly slower than running all training data through the network in inference mode, because the backbone is frozen. The slowest part is the inversions of the Gram matrix, during selection of $\lambda$, but even for $M=10000$, this is in the order of 1 minute per task on a CPU, which can be easily optimized further if needed. Our view is that the efficiency and simplicity of our approach compared to the alternatives is very strong.
C.4 Compute
@@ F.3 Comparison of impact of Phase 1 for different numbers of CIL tasks
Therefore, in this section we show in Table A4 results for RanPAC only for the case where variability has greater effect, i.e. when Phase 1 is included, and AdaptMLP chosen. VTAB is excluded from this analysis, since it is a special case where it makes sense to consider only the case of $T=5$. The comparison results for L2P, DualPrompt and ADaM are copied from zhou2023revisiting.
@@
Table A4: Comparison of CIL results for different number of tasks. For this table, the AdaptMLP PETL method was used for RanPAC. The comparison results for L2P, DualPrompt and ADaM are copied from zhou2023revisiting; for ADaM, the best performing PETL method was used. In most cases, Average Accuracy decreases as $T$ increases, with $T=20$ typically not significantly better than the case of no PETL (“No Phase 1”). Values marked with $*{}^{*}$ indicates data is for $T=6$ tasks, instead of $T=5$. Note that accuracies for comparison methods for Cars were not available from zhou2023revisiting
@@ F.4 Task Agnostic Continual Learning
Unlike CIL and DIL, ‘task agnostic’ CL is a scenario where there is no clear concept of a ‘task’ during training zeno2019task. The concept is also known as ‘task-free’ Shanahan21. It contrasts with standard CIL where although inference is task agnostic, training is applied to disjoint sets of classes, described as tasks. To illustrate the flexibility of RanPAC, we show here that it is simple to apply it to task agnostic continual learning. We use the Gaussian scheduled CIFAR100 protocol of L2P, which was adapted from Shanahan21. We use 200 ‘micro-tasks’ which sample from a gradually shifting subset of classes, with 5 batches of 48 samples in each micro-task. There are different possible choices for how to apply Algorithm 1. For instance, the ‘first session’ for Phase 1 could be defined as a particular total number of samples trained on, e.g. 10% of the anticipated total number of samples. Then in Phase 2, the outer for loop over tasks could be replaced by a loop over all batches, or removed entirely. In both cases, the result for ${\bf G}$ and ${\bf C}$ will be unaffected. The greater challenge is in determining $\lambda$, but generally for a large number of samples, $\lambda$ can be small, or zero. For a small number of training samples, a queue of samples could be maintained such that the oldest sample in the queue is used to update ${\bf G}$ and ${\bf C}$, with all newer samples used to compute $\lambda$ if inference is required, and then all samples in the buffer added to ${\bf G}$ and ${\bf C}$.